Similar Posts
Optical Fiber Cable Rejoin With Poll
Posted onDue To Strom the optical fiber cable fall down near DTW-114 to 117 and chaira brize to DTW-218 now everything is normal
ACTIVITY | DAILY REPORTS | Parvez | REPORTSOGR Booster room dewatering pump run command showing problem at low level
Posted onFloat switch reconnection and rearrange. Now normal
ACTIVITY | DAILY REPORTS | DTWDTW-113 Scada Pack Install
Posted onDTW-113 Scada Pack Programing And Install Done
ACTIVITY | DAILY REPORTS | DTW | PappuDTW – 111 LIT Maintenance Work Complete.
Posted onDtw 111 LIT Reinstallation complete . DTW 111 থেকে পুরাতন একটি LIT স্টোরে জমা দেয়া হয়েছে ৷
IRP-2 Server Rack Installation and Optical Fiber Rejoin
Posted onIRP-2 Server Rack Installation and Optical Fiber Rejoin Done
ACTIVITY | DAILY REPORTS | Jewel | Pappu | ParvezOGR booster pump D winding temperature & Bearing temperature sensor checking.
Posted onProblem: In those sensors, high temperature shows in transmitter display. But when we check the mA current of sensor signal, that was OK….
